Given the following demand equation P = 60- 2Q and the cost equation TC(q) = 100+3Q2

For the market of shoes

a.If the enterprises in this market take advantages of perfectly competitive market, find the equilibrium outputs and price.

b.If the seller in this market takes the advantages of pure monopoly, find the equilibrium outputs and price.

c.If the seller in this market takes the advantages ofoligopolistic market, find the equilibrium outputs and price.


1
Expert's answer
2022-03-21T12:54:03-0400

For perfectly competitive markets,

the enterprises are price takers


P= 60-2Q

TC= 100+3Q"^2"

For price takers

P = MC =MR

MC = 6Q

In equilibrium,

MC= P

6Q= 60-2Q

Q=7.5

While P is,

P= 60- 2(7.5)

P= 60-15

P= 45

b) For a monopoly equilibrium is;

MC = MR

Total revenue

P×Q = 60Q-2Q"^2"

MR = 60- 4Q

In equilibrium,

60-4Q = 6Q

Q = 6

Price,

"P= 60 - 2(6)"

"P = 48"

c) For an Oligopoly equilibrium is the same as in Monopolistic market;

MC = MR

Total revenue

P×Q = 60Q-2Q"^2"

MR = 60- 4Q

In equilibrium,

60-4Q = 6Q

Q = 6

Price,

"P= 60 - 2(6)"

"P = 48"